OpenStudy (anonymous):

Barry is trying round √643 to the nearest tenth. This is his work: 252 = 625 ; 262 = 676 25.32 = 640.09 25.352 = 642.6225 25.362 = 643.1296 Barry chose 25.36. What did Barry do wrong?

OpenStudy (anonymous):

\[25^2 = 625 ; 26^2 = 676 25.3^2 = 640.09 25.35^2 = 642.6225 25.36^2 = 643.1296\]